OpenStudy (anonymous):

How are these concepts of direct, inverse, and joint variation used in everyday life? Provide examples for each.

OpenStudy (anonymous):

direct: \[y \alpha x \rightarrow y = kx\]=\]

OpenStudy (anonymous):

inverse: \[y \alpha \frac{ 1 }{ x }\rightarrow y = \frac{ k }{ x }\]

OpenStudy (anonymous):

y varies jointly as x and a constant

OpenStudy (anonymous):

jointly: \[y = mx + c\]

OpenStudy (anonymous):

The first is y varies directly as x The second is y varies inversely as x The third is y varies jointly as x and a constant where k is the proportionality constant